OpenStudy (vera_ewing):

Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created?

OpenStudy (misty1212):

HI again!!

OpenStudy (misty1212):

you need to know two things, a) what it looks like and b) the vertex

OpenStudy (vera_ewing):

f(x) = 1/16(x - 1)^2 + 2

OpenStudy (vera_ewing):

Is that the answer ^^ can you please check?

OpenStudy (misty1212):



OpenStudy (misty1212):

vertex is \((1,2)\)

OpenStudy (misty1212):

oh yeah your answer is right

OpenStudy (vera_ewing):

Oh ok. Thanks! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

\[4p(y-k)=(x-h)^2\]

OpenStudy (misty1212):

\[16(y-2)=(x-1)^2\]